# Information geometry of Lévy processes and financial models

### v2

@@ Abstract.
We explore the information geometry of Lévy processes. As a starting point, we derive the $\alpha$-divergence between two Lévy processes. Subsequently, the Fisher information matrix and the $\alpha$-connection associated with the geometry of Lévy processes are computed from the $\alpha$-divergence. In addition, we discuss statistical applications of this information geometry. As illustrative examples, we investigate the differential-geometric structures of various Lévy processes relevant to financial modeling, including tempered stable processes, the CGMY model, and variance gamma processes.
@@ 1. Introduction
Information geometry, which studies the differential-geometric structure of statistical models, has emerged as a powerful tool for understanding complex probabilistic systems. By representing families of probability distributions as Riemannian manifolds, this framework enables a geometric interpretation of statistical inference and parameter estimation procedures [15, 16, 17, 1, 25, 19]. The metric tensors and connections on statistical manifolds are derived from divergence functions, in particular, the $\alpha$-divergence, a generalization of the Kullback–Leibler divergence. In this context, the metric tensor corresponds to the Fisher information matrix, which plays a fundamental role in statistical theory and applications. In recent decades, information geometry has been successfully applied to exponential families [15, 16, 17, 26, 27, 1, 19], time series models [32, 31, 25, 35, 3, 4, 34, 12, 11, 29, 7], Bayesian prediction [25, 35, 34, 12, 11, 29], and, more recently, to stochastic processes with non-Gaussian features [8].
@@
In previous work [8], the information geometry of tempered stable processes, a subclass of Lévy processes characterized by heavy tails and jumps, was investigated. The $\alpha$-divergence between two tempered stable processes was derived as a two-fold generalization of the result by Kim and Lee [21], in which the Kullback–Leibler divergence (relative entropy) for classical tempered stable processes, a subset of tempered stable processes, was obtained. Moreover, the corresponding Fisher information matrix and $\alpha$-connection, which capture the geometric structure of these processes, were constructed from the $\alpha$-divergence. Several statistical applications from the information geometry of tempered stable processes were also provided.
@@
The differential-geometric approach to tempered stable processes opens new horizons for its applications in finance. Due to their ability to capture characteristic features of asset returns such as asymmetry and fat-tails, various tempered stable processes have been widely employed in financial modeling, as the normality assumption is empirically invalid in this domain. Their use in finance spans a range of models, including time series analysis via ARMA-GARCH models with tempered stable distributed innovations [22, 23, 24], as well as applications in portfolio management, risk management, and option pricing [36, 5, 10, 9, 18, 2, 20].
@@
However, tempered stable processes, while rich and flexible for financial modeling, constitute only a subset of the broader class of Lévy processes. Lévy processes include not only various tempered stable processes but also stable, compound Poisson, variance gamma processes, and other models [14, 30, 28]. These processes are characterized by stationary and independent increments and serve as fundamental models in the study of jump processes and infinite divisibility. Extending the information geometry developed for tempered stable processes [8] to the full class of Lévy processes is a natural and necessary progression, with implications for both theoretical development and practical applications.
@@
The present paper aims to undertake this extension. Our main goal is to develop the information geometry of Lévy processes. As mentioned above, we generalize the construction of information-geometric structures from the tempered stable process setting [8] to the full class of Lévy processes. This extension also provides an alternative generalization of the result in Cont and Tankov [14], where the Kullback–Leibler divergence for Lévy processes was derived. We begin by deriving the $\alpha$-divergence between two Lévy processes. Since divergence functions assess dissimilarity between probability distributions or statistical models, they underlie the definition of key geometric objects such as the Fisher information matrix and the $\alpha$-connection, derived from the $\alpha$-divergence [1, 13]. We also discuss potential statistical applications for Lévy processes from the geometric perspective.
@@
Having established the general framework, we investigate several concrete examples of Lévy processes relevant to financial modeling. These include tempered stable processes [30], CGMY models [6], and variance gamma processes [28], all of which have been widely used to capture the stylized features of financial asset returns. For each model, we derive the $\alpha$-divergence and construct the associated information geometry such as Fisher information matrix and $\alpha$-connection. Moreover, we provide the statistical advantages originated from the geometry of each model.
@@
By developing a unified geometric viewpoint on Lévy processes, this work contributes to the growing intersection of stochastic processes, mathematical finance, and statistics through the frame of information geometry. Our results offer both theoretical insights and practical tools for analyzing Lévy processes and jump-driven models. In doing so, we build a bridge between abstract geometric concepts and concrete applications, particularly in domains such as finance where Lévy processes play a central role.
@@
This paper is organized as follows. In Section 2, we revisit the fundamentals of Lévy processes and review relevant prior work. Section 3 develops the information geometry of Lévy processes, including $\alpha$-divergence, Fisher information matrix, and $\alpha$-connection, and also discusses its statistical applications. In Section 4, we present several illustrative examples with financial applications. Finally, we conclude the paper with a summary.
@@ 2. Lévy processes
Lévy processes constitute a broad and fundamental class of stochastic processes characterized by stationary and independent increments. By generalizing both Brownian motion and Poisson processes, they serve as canonical building blocks for modeling random phenomena that exhibit discontinuities, jumps, and heavy tails. This versatility makes Lévy processes particularly suitable for describing systems influenced by random shocks and bursts. They are especially important in fields such as quantitative finance, physics, signal processing, and biology, where non-Gaussian and discontinuous behavior is frequently observed.
@@
The Lévy–Itô decomposition states that a Lévy process can be uniquely expressed as the combination of a Brownian (diffusion) component, a drift term, and a pure jump component governed by a Lévy measure. This decomposition for a Lévy process is fully characterized by its characteristic (Lévy) triplet $(\sigma,\nu,\gamma)$, where $\sigma\geq 0$ is the diffusion coefficient, $\nu$ is a Lévy measure on $\mathbb{R}$, and $\gamma\in\mathbb{R}$ is the drift term. The Lévy measure $\nu$ is a positive measure satisfying:
@@
A cornerstone of Lévy process theory is the Lévy–Khintchine formula, which characterizes Lévy processes via their characteristic functions [30]. According to the Lévy–Khintchine formula, the characteristic function of a Lévy process $(X_{t},\mathbb{P})_{t\in[0,T]}$ with Lévy triplet $(\sigma,\nu,\gamma)$ is given by
@@
where $(U_{t},\mathbb{U})_{t\in[0,T]}$ is a Lévy process with characteristic triplet $(\sigma_{\mathbb{U}},\nu_{\mathbb{U}},\gamma_{\mathbb{U}})$ given by
@@
In brief, the proof relies on applying the Radon–Nikodym derivative from Proposition 1 to the definition of the Kullback–Leibler divergence. We will revisit this proof in the following section.
@@ 3. Information geometry of Lévy processes
In this section, we explore the information geometry of Lévy processes. We begin by deriving the $\alpha$-divergence of Lévy processes from the definition of $f$-divergence. We then investigate the associated geometric structures, including Fisher information matrix and $\alpha$-connection, and discuss statistical applications.
@@
(17) $f(t)=t\log{t}-t+1.$
Another important example of $f$-divergence is the $\alpha$-divergence, which plays a central role in information geometry [1]. The $\alpha$-divergence can be expressed as the $f$-divergence corresponding to the following choice of $f$:
@@
It is straightforward to verify that the functions $f^{(\alpha)}$ for $\alpha=\pm 1$ can be obtained by applying L’Hôpital’s rule to the $f$-divergence for $\alpha\neq\pm 1$ and taking the limit as $\alpha\to\pm 1$, respectively. Moreover, the $\alpha=-1$ case in Eq. (21) coincides with Eq. (17), therefore recovering the Kullback–Leibler divergence.
@@
Based on the definition of the Kullback–Leibler divergence given by Eq. (27) together with Eq. (17), Cont and Tankov [14] computed the Kullback–Leibler divergence for Lévy processes under Proposition 1 by evaluating the following expression:
(28) $KL(\mathbb{P}||\mathbb{Q})=\mathbb{E}_{\mathbb{Q}}[f({\rm{e}}^{U_{T}})],$
where $f$ is given by Eq. (17) and $U_{T}$ is defined in Proposition 1. This calculation consists of the basis for the proof of Proposition 2, as mentioned earlier.
Analogous to Proposition 2, the $\alpha$-divergence between Lévy processes can be derived by applying Eq. (21) to Eq. (27):
@@
where $\Phi_{T}(z;\boldsymbol{\xi})$ is given by Eq. (3) and $U_{T}$ is defined in Proposition 1.
@@
When $\sigma\neq 0$, $\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})$ is obtained as
$T^{-1}\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})=\Big(\frac{1-\alpha}{2}\Big)^{2}\frac{\sigma_{\mathbb{U}}^{2}}{2}+\frac{1-\alpha}{2}\gamma_{\mathbb{U}}+\int_{-\infty}^{\infty}({\rm e}^{\frac{1-\alpha}{2}y}-1-\frac{1-\alpha}{2}y\mathbbm{1}_{|y|\leq 1})\nu_{\mathbb{U}}(dy).$
By applying Eq. (9), Eq. (10), Eq. (11), and Eq. (12), $T^{-1}\Phi_{T}\left(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}}\right)$ can be expressed as
@@
$=\Big(\frac{1-\alpha}{2}\Big)^{2}\frac{(\sigma\eta)^{2}}{2}+\frac{1-\alpha}{2}\Big(-\frac{(\sigma\eta)^{2}}{2}-\int_{-\infty}^{\infty}({\rm e}^{y}-1-y\mathbbm{1}_{|y|\leq 1})(\nu_{\mathbb{Q}}\psi^{-1})(dy)\Big)$
 $\quad\textrm{ }+\int_{-\infty}^{\infty}({\rm e}^{\frac{1-\alpha}{2}y}-1-\frac{1-\alpha}{2}y\mathbbm{1}_{|y|\leq 1})(\nu_{\mathbb{Q}}\psi^{-1})(dy)$
$=-\frac{1-\alpha^{2}}{4}\frac{(\sigma\eta)^{2}}{2}-\int_{-\infty}^{\infty}(\frac{1-\alpha}{2}{\rm e}^{y}+\frac{1+\alpha}{2}-{\rm e}^{\frac{1-\alpha}{2}y})(\nu_{\mathbb{Q}}\psi^{-1})(dy)$
$=-\frac{1-\alpha^{2}}{4}\frac{(\sigma\eta)^{2}}{2}-\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx)$
@@
From the equation above, we derive Eq. (31).
@@
Using this term, it is straightforward to verify that substituting the expression above into Eq. (31) yields Eq. (32).
When $\sigma=0$, $\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})$ is obtained as
$T^{-1}\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})$ $=\int_{-\infty}^{\infty}({\rm e}^{\frac{1-\alpha}{2}y}-1-\frac{1-\alpha}{2}y\mathbbm{1}_{|y|\leq 1})\nu_{\mathbb{U}}(dy)$
$=-\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx).$
@@
By using the logarithmic characteristic function $\Phi_{t}(z;\boldsymbol{\xi})$ from Eq. (3), the $\alpha$-divergence for $\alpha\neq\pm 1$ can be expressed as
@@
When $\alpha=-1$, the Kullback–Leibler divergence can be obtained in two different ways. The first approach is a direct calculation by plugging the $\alpha=-1$ case in Eq. (21) into Eq. (29), which yields the same result with the derivation presented by Cont and Tankov [14]. The second method involves applying L’Hôpital’s rule to the expression for $\alpha$-divergence with $\alpha\neq\pm 1$ and taking the limit as $\alpha\to-1$, which also recovers the Kullback–Leibler divergence.
@@
Similarly, there are two ways to obtain the $1$-divergence: one by leveraging L’Hôpital’s rule, and the other by direct calculation from the definition of $\alpha$-divergence. Additionally, a third approach involves using $\alpha$-duality to derive the dual Kullback–Leibler divergence:
@@
We now return to the case of $\sigma=0$. As noted earlier, when $\sigma=0$, the first terms in Eq. (31) and Eq. (32) vanish by Proposition 3, allowing Eq. (51) to be derived directly from Eq. (40) and Eq. (47):
@@
It is worth noting that the Kullback–Leibler divergences in Eq. (40) and Eq. (47), i.e., the cases with $\alpha=-1$, are identical to those presented in Proposition 2, as derived by Cont and Tankov [14].
@@
(57) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big(1-{\rm{e}}^{-\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}\Big).&(\alpha\neq\pm 1)\\ \lim_{\alpha\to-1}-\frac{2}{\alpha}\frac{\partial_{\alpha}\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}{\partial\alpha}&(\alpha=-1)\\ \lim_{\alpha\to 1}-\frac{2}{\alpha}\frac{\partial_{\alpha}\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}{\partial\alpha}&(\alpha=1)\end{array}\right.$.
@@
The proof is straightforward. For $\alpha\neq\pm 1$, the result follows directly from Eq. (53). When $\alpha=\pm 1$, as discussed in the proof of Theorem 1, we apply L’Hôpital’s rule to Eq. (53) and take the limit as $\alpha\to\pm 1$:
@@
Let $(X_{t},\mathbb{P})_{t\in[0,T]}$ be a Lévy process on $(\Omega,\mathcal{F})$ with the Lévy triplet of $(\sigma,\nu,\gamma)$. When $\sigma\neq 0$, the Fisher information matrix of Lévy processes is represented with
@@
and the $\alpha$-connection for information geometry of Lévy processes is given as
@@
and the $\alpha$-connection for information geometry of Lévy processes is given as
@@
As done in Theorem 1, we divide the analysis into two cases: $\sigma\neq 0$ and $\sigma=0$.
We begin with the case of $\sigma\neq 0$. By substituting Eq. (40) into Eq. (66) and Eq. (67), we obtain Eq. (60) and Eq. (61).
@@
In the case of $\sigma=0$, the first terms of the exponents in Eq. (40) and Eq. (47) vanish. Consequently, the first terms in Eq. (60) and Eq. (61) (or equivalently, in Eq. (62) and Eq. (63)) also become zero. A direct computation using Eq. (51) in Eq. (66) and Eq. (67) yields the same geometric structures. ∎
@@
Let $(X_{t},\mathbb{P})_{t\in[0,T]}$ be a Lévy process on $(\Omega,\mathcal{F})$ with the Lévy triplet of $(\sigma,\nu,\gamma)$. The Fisher information matrix of Lévy processes and the $\alpha$-connection for information geometry of Lévy processes are given as
@@
(70) $g_{ij}$ $=-\frac{4}{1-\alpha^{2}}\Delta^{(\alpha)}_{T}(\partial_{i},\tilde{\partial}_{j}){\rm e}^{-\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}},$
(71) $\Gamma_{ij,k}^{(\alpha)}$ $=-\frac{4}{1-\alpha^{2}}\Delta^{(\alpha)}_{T}(\partial_{i}\partial_{j},\tilde{\partial}_{k}){\rm e}^{-\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}}.$
@@
For instance, when the information geometry is e-flat, it is possible to employ the penalized log-likelihood approach [17]:

### v3

@@ Abstract.
We develop the information geometry of Lévy processes. Deriving $\alpha$-divergences directly in terms of the Lévy triplets of the Lévy processes, we identify Fisher information matrix and $\alpha$-connection on the statistical manifold. In addition, we discuss statistical implications of this information geometry, including bias reduction estimation and Bayesian predictive priors. Several Lévy processes, broadly used for financial modeling such as tempered stable processes, the CGMY model, variance gamma processes, and the Merton model, are investigated through their differential-geometric structures as illustrative examples.
@@ 1. Introduction
Information geometry, which studies the differential-geometric structure of statistical models, has emerged as a powerful tool for understanding complex probabilistic systems. By representing families of probability distributions as Riemannian manifolds, this framework enables a geometric interpretation of statistical inference and parameter estimation procedures [15, 16, 17, 1, 25, 19]. The geometric structures of statistical manifolds such as metric tensors and connections are derived from divergence functions, in particular, the $\alpha$-divergence that is a generalization of the Kullback–Leibler divergence. In this context, the metric tensor corresponds to the Fisher information matrix, which plays a fundamental role in statistical theory and applications. In recent decades, information geometry has been successfully applied to exponential families [15, 16, 17, 26, 27, 1, 19], time series models [33, 32, 25, 36, 3, 4, 35, 12, 11, 30, 7], Bayesian prediction [25, 36, 35, 12, 11, 30], and more recently to stochastic processes with non-Gaussian features [8].
@@
In previous work [8], tempered stable processes, a subclass of Lévy processes characterized by heavy tails and jumps, was investigated from the perspective of information geometry. The $\alpha$-divergence between two tempered stable processes was derived as a two-fold generalization of Kim and Lee [21] where the Kullback–Leibler divergence of classical tempered stable processes, a subset of tempered stable processes, was obtained. Moreover, the corresponding Fisher information matrix and $\alpha$-connection, which capture the geometric structure of these processes, were constructed from the $\alpha$-divergence. Several statistical applications from the information geometry of tempered stable processes were also provided.
@@
The differential-geometric approach to tempered stable processes opens new horizons for its applications in finance. Due to their ability to describe characteristic features of asset returns such as asymmetry and fat-tails, various tempered stable processes have been widely employed in financial modeling, where the normality assumption is empirically invalid. Their use in finance spans a range of models, including time series analysis via ARMA-GARCH models with tempered stable distributed innovations [22, 23, 24], as well as applications in portfolio management, risk management, and option pricing [37, 5, 10, 9, 18, 2, 20].
@@
However, tempered stable processes, while rich and flexible for financial modeling, constitute only a subset of the broader class of Lévy processes. Lévy processes include not only various tempered stable processes but also stable processes, compound Poisson processes, variance gamma processes, and other models [14, 31, 28]. These processes serve as fundamental models in the study of jump processes and infinite divisibility. Extending the information geometry developed for tempered stable processes [8] to the full class of Lévy processes is a natural and necessary progression with implications for both theoretical development and practical applications.
@@
The present paper aims to provide this extension. Our main goal is to develop the information geometry of Lévy processes. As mentioned above, we generalize the construction of information-geometric structures from the tempered stable process setting [8] to the full class of Lévy processes. This extension also provides an alternative generalization of the result in Cont and Tankov [14], where the Kullback–Leibler divergence for Lévy processes was derived. We begin by deriving the $\alpha$-divergence between two Lévy processes. Since divergence functions evaluate the dissimilarity between probability distributions or statistical models, they underlie the definition of key geometric objects such as Fisher information matrix and $\alpha$-connection, derived from the $\alpha$-divergence [1, 13]. We also discuss potential statistical applications for Lévy processes from the geometric perspective.
@@
Having established the general framework, we investigate several concrete examples of Lévy processes relevant to financial modeling. These include tempered stable processes [31], the CGMY model [6], variance gamma processes [28], and the Merton model [29], all of which have been widely used to model the stylized features of financial asset returns. For each model, we derive the $\alpha$-divergence and construct the associated information geometry such as Fisher information matrix and $\alpha$-connection. Moreover, the statistical advantages originated from the statistical manifolds of these processes are discussed.
@@
By developing a unified geometric viewpoint on Lévy processes, this work contributes to the growing intersection of stochastic processes, mathematical finance, and statistics within the framework of information geometry. Our results offer both theoretical insights and practical tools for analyzing Lévy processes and financial models. Leveraging these advantages, we build a bridge between abstract geometric concepts and concrete applications in domains such as finance where Lévy processes play a central role.
@@
This paper is organized as follows. In Section 2, we revisit the fundamentals of Lévy processes and the relevant prior work. Section 3 develops the information geometry of Lévy processes, including $\alpha$-divergence, Fisher information matrix, and $\alpha$-connection, and also discusses its statistical applications. In Section 4, we present several illustrative examples with financial applications. Finally, we conclude the paper.
@@ 2. Lévy processes
Lévy processes constitute a broad and fundamental class of stochastic processes characterized by stationary and independent increments. By generalizing both Brownian motion and Poisson processes, the processes become canonical building blocks for modeling random phenomena that exhibit discontinuities, jumps, and heavy tails. This versatile aspect makes Lévy processes particularly suitable for describing systems influenced by random shocks and bursts. They are especially important in fields such as finance, physics, signal processing, and biology, where non-Gaussian and discontinuous behavior is frequently observed.
@@
The Lévy–Itô decomposition states that a Lévy process can be uniquely expressed as the combination of a Brownian (diffusion) component, a drift term, and a pure jump component governed by a Lévy measure. This decomposition for a Lévy process is fully described by its characteristic (Lévy) triplet $(\sigma,\nu,\gamma)$, where $\sigma\geq 0$ is the diffusion coefficient, $\nu$ is the Lévy measure on $\mathbb{R}$, and $\gamma\in\mathbb{R}$ is the drift term. The Lévy measure $\nu$ is a positive measure satisfying the following conditions:
@@
A cornerstone of Lévy process theory is the Lévy–Khintchine formula, which characterizes Lévy processes via their characteristic functions [34, 31]. According to the Lévy–Khintchine formula, the characteristic function of a Lévy process $(X_{t},\mathbb{P})_{t\in[0,T]}$ with a Lévy triplet $(\sigma,\nu,\gamma)$ is given by
@@
where $(U_{t},\mathbb{U})_{t\in[0,T]}$ is a Lévy process with a characteristic triplet $(\sigma_{\mathbb{U}},\nu_{\mathbb{U}},\gamma_{\mathbb{U}})$ given by
@@
In brief, the proof of Proposition 2 relies on applying the Radon–Nikodym derivative of Proposition 1 to the definition of the Kullback–Leibler divergence. We will revisit this proof in the following section.
@@ 3. Information geometry of Lévy processes
In this section, we explore the information geometry of Lévy processes. Starting with deriving the $\alpha$-divergence of Lévy processes from the definition of $f$-divergence, we investigate the associated geometric structures, including Fisher information matrix and $\alpha$-connection. The discussion of its statistical applications follows.
@@
(17) $f_{KL}(t)=t\log{t}-t+1.$
Another important example of $f$-divergence is the $\alpha$-divergence, which plays a central role in information geometry [1]. The $\alpha$-divergence can be expressed as the $f$-divergence with the following choice of $f$:
@@
It is straightforward to verify that the functions $f^{(\alpha)}$ for $\alpha=\pm 1$ can be obtained by applying L’Hôpital’s rule to the $f$-divergence for $\alpha\neq\pm 1$ and taking the limit of $\alpha\to\pm 1$, respectively. Moreover, the $\alpha=-1$ case in Eq. (21) coincides with Eq. (17) for the Kullback–Leibler divergence.
@@
Based on Eq. (27), Cont and Tankov [14] computed the Kullback–Leibler divergence for Lévy processes under Proposition 1 by evaluating the following expression:
(28) $KL(\mathbb{P}||\mathbb{Q})=\mathbb{E}_{\mathbb{Q}}[f_{KL}({\rm{e}}^{U_{T}})],$
where $f_{KL}$ is given by Eq. (17) and $U_{T}$ is defined in Proposition 1. This calculation consists of the basis for the proof of Proposition 2, as mentioned earlier.
Analogous to Proposition 2, the $\alpha$-divergence between Lévy processes is derived by applying Eq. (21) to Eq. (27):
@@
where $\Phi_{T}(z;\boldsymbol{\xi})$ is given by Eq. (3) and $\boldsymbol{\xi}_{\mathbb{U}}$ is the parameters of $U_{T}$ defined in Proposition 1.
@@
By applying Eq. (9), Eq. (10), and Eq. (11), $T^{-1}\Phi_{T}\left(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}}\right)$ can be expressed with
(34) $\begin{aligned} &T^{-1}\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})\\ &=\Big(\frac{1-\alpha}{2}\Big)^{2}\frac{\sigma_{\mathbb{U}}^{2}}{2}+\frac{1-\alpha}{2}\gamma_{\mathbb{U}}+\int_{-\infty}^{\infty}({\rm e}^{\frac{1-\alpha}{2}y}-1-\frac{1-\alpha}{2}y\mathbbm{1}_{|y|\leq 1})\nu_{\mathbb{U}}(dy)\\ &=\Big(\frac{1-\alpha}{2}\Big)^{2}\frac{(\sigma\eta)^{2}}{2}+\frac{1-\alpha}{2}\Big(-\frac{(\sigma\eta)^{2}}{2}-\int_{-\infty}^{\infty}({\rm e}^{y}-1-y\mathbbm{1}_{|y|\leq 1})(\nu_{\mathbb{Q}}\psi^{-1})(dy)\Big)\\ &\quad\textrm{ }+\int_{-\infty}^{\infty}({\rm e}^{\frac{1-\alpha}{2}y}-1-\frac{1-\alpha}{2}y\mathbbm{1}_{|y|\leq 1})(\nu_{\mathbb{Q}}\psi^{-1})(dy)\\ &=-\frac{1-\alpha^{2}}{4}\frac{(\sigma\eta)^{2}}{2}-\int_{-\infty}^{\infty}(\frac{1-\alpha}{2}{\rm e}^{y}+\frac{1+\alpha}{2}-{\rm e}^{\frac{1-\alpha}{2}y})(\nu_{\mathbb{Q}}\psi^{-1})(dy)\\ &=-\frac{1-\alpha^{2}}{4}\frac{(\sigma\eta)^{2}}{2}-\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx).\end{aligned}$
When $\sigma\neq 0$, plugging Eq. (12) to Eq. (34) yields $T^{-1}\Phi_{T}\left(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}}\right)$ as
@@
From the equation above, we obtain Eq. (31).
@@
It is straightforward to verify that substituting the expression above into Eq. (31) yields Eq. (32).
When $\sigma=0$ in Eq. (34), $\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})$ is obtained as
$T^{-1}\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})$ $=-\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx).$
@@
When $\sigma=0$, the $\alpha$-divergence between two Lévy processes are given by
@@
By using the logarithmic characteristic function $\Phi_{t}(z;\boldsymbol{\xi})$ from Eq. (3), the $\alpha$-divergence for $\alpha\neq\pm 1$ can be expressed with
@@
When $\alpha=-1$, the Kullback–Leibler divergence can be obtained in two different ways. The first approach is a direct calculation by plugging the $\alpha=-1$ case in Eq. (21) into Eq. (29), which yields the same result with the derivation presented by Cont and Tankov [14]. The second method involves applying L’Hôpital’s rule to the expression for $\alpha$-divergence with $\alpha\neq\pm 1$ and taking the limit of $\alpha\to-1$, which also recovers the Kullback–Leibler divergence.
@@
Similarly, there are two ways to obtain the $1$-divergence: one by leveraging L’Hôpital’s rule, and the other by direct calculation from the definition of $\alpha$-divergence. Besides, the third approach involves using $\alpha$-duality to derive the dual Kullback–Leibler divergence:
@@
We now return to the case of $\sigma=0$. As noted earlier, when $\sigma=0$, the first terms in Eq. (31) and Eq. (32) vanish by Proposition 3 due to Eq. (7) and Eq. (12), allowing Eq. (52) to be derived directly from Eq. (41) and Eq. (48):
@@
It is straightforward to check that the Kullback–Leibler divergences in Eq. (41) and Eq. (48), i.e., the cases with $\alpha=-1$, are identical to those presented in Proposition 2, as derived by Cont and Tankov [14].
@@
(58) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big(1-{\rm{e}}^{-\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}\Big)&(\alpha\neq\pm 1)\\ \lim_{\alpha\to-1}-\frac{2}{\alpha}\frac{\partial_{\alpha}\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}{\partial\alpha}&(\alpha=-1)\\ \lim_{\alpha\to 1}-\frac{2}{\alpha}\frac{\partial_{\alpha}\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}{\partial\alpha}&(\alpha=1)\end{array}\right.$.
@@
The proof is straightforward. For $\alpha\neq\pm 1$, the result follows directly from Eq. (54). When $\alpha=\pm 1$, as discussed in the proof of Theorem 1, we apply L’Hôpital’s rule to Eq. (54) and take the limit of $\alpha\to\pm 1$:
@@
Let $(X_{t},\mathbb{P})_{t\in[0,T]}$ be a Lévy process on $(\Omega,\mathcal{F})$ with a Lévy triplet $(\sigma,\nu,\gamma)$. When $\sigma\neq 0$, the Fisher information matrix of Lévy processes is represented with
@@
and the $\alpha$-connection for information geometry of Lévy processes is given by
@@
and the $\alpha$-connection for information geometry of Lévy processes is given by
@@
As done in Theorem 1, we divide the proof into two cases: $\sigma\neq 0$ and $\sigma=0$.
We begin with the case of $\sigma\neq 0$. By substituting Eq. (41) into Eq. (67) and Eq. (68), we obtain Eq. (61) and Eq. (62), respectively.
@@
In the case of $\sigma=0$, the first terms of the exponents in Eq. (41) and Eq. (48) vanish. Consequently, the first terms in Eq. (61) and Eq. (62) (or equivalently, in Eq. (63) and Eq. (64)) also become zero. A direct computation using Eq. (52) in Eq. (67) and Eq. (68) yields the same geometric structures of Eq. (65) and Eq. (66), respectively. ∎
@@
Let $(X_{t},\mathbb{P})_{t\in[0,T]}$ be a Lévy process on $(\Omega,\mathcal{F})$ with the Lévy triplet of $(\sigma,\nu,\gamma)$. The Fisher information matrix of Lévy processes and the $\alpha$-connection for information geometry of Lévy processes are given by
@@
(71) $g_{ij}$ $=-\frac{4}{1-\alpha^{2}}\Delta^{(\alpha)}_{T}(\partial_{i}||\tilde{\partial}_{j}){\rm e}^{-\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}},$
(72) $\Gamma_{ij,k}^{(\alpha)}$ $=-\frac{4}{1-\alpha^{2}}\Delta^{(\alpha)}_{T}(\partial_{i}\partial_{j}||\tilde{\partial}_{k}){\rm e}^{-\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}}.$
@@
For instance, when the information geometry is e-flat where the connection is vanishing with $\alpha=1$, it is possible to employ the penalized log-likelihood approach [17]: